# Unifying Search and Recommendation: A Generative Paradigm Inspired by Information Theory

### v2

@@ Abstract.
Recommender systems and search engines serve as foundational elements of online platforms, with the former delivering information proactively and the latter enabling users to seek information actively. Unifying both tasks in a shared model is promising since it can enhance user modeling and item understanding. Previous approaches mainly follow a discriminative paradigm, utilizing shared encoders to process input features and task-specific heads to perform each task. However, this paradigm encounters two key challenges: gradient conflict and manual design complexity. From the information theory perspective, these challenges potentially both stem from the same issue — low mutual information between the input features and task-specific outputs during the optimization process.
To tackle these issues, we propose GenSR, a novel generative paradigm for unifying search and recommendation (S&R), which leverages task-specific prompts to partition the model’s parameter space into subspaces, thereby enhancing mutual information. To construct effective subspaces for each task, GenSR first prepares informative representations for each subspace and then optimizes both subspaces in one unified model. Specifically, GenSR consists of two main modules: (1) Dual Representation Learning, which independently models collaborative and semantic historical information to derive expressive item representations; and (2) S&R Task Unifying, which utilizes contrastive learning together with instruction tuning to generate task-specific outputs effectively. Extensive experiments on two public datasets show GenSR outperforms state-of-the-art methods across S&R tasks. Our work introduces a new generative paradigm compared with previous discriminative methods and establishes its superiority from the mutual information perspective.
@@ 1. Introduction
Recommender systems and search engines have become indispensable components of modern online service platforms, including e-commerce websites and social media networks (Zhang et al., 2024; Wu et al., 2024). Usually, search and recommendation (S&R) tasks are trained on separate models (Bhattacharya et al., 2024). However, unifying these two tasks within a shared model offers significant advantages (Yao et al., 2021; Si et al., 2023): First, joint modeling of user behaviors across both S&R provides a more comprehensive understanding of user preferences. Second, leveraging interactions from both task domains enriches item-side interactions, enhancing the understanding and delivery of items. Therefore, it is promising to develop a unified model for S&R.
@@
Figure 1. (a) and (b) illustrate the discriminative and generative paradigms for unifying S&R, respectively, highlighting their differences in parameter space. While the discriminative paradigm results in low mutual information by optimizing S&R tasks within a shared parameter space, the generative paradigm enhances mutual information by partitioning the space into distinct subspaces through prompts.
@@
Existing efforts to unify S&R mostly follow a discriminative paradigm, which utilizes shared encoders to process input features (e.g., user interaction history) and task-specific heads to predict whether a user will click on an item in S&R tasks, as illustrated in Figure 1(a) (Shi et al., 2024; Xie et al., 2024). However, such paradigm has two primary drawbacks: 1) Gradient conflict: The optimization of shared parameters may introduce gradient conflict, thereby limiting the model’s performance (Shi et al., 2023; He et al., 2024). 2) Manual design complexity: Determining the extent of parameter sharing and designing task-specific heads requires significant manual effort, increasing complexity and limiting extensibility (Rahimian et al., 2023).
@@
Inspired by information theory, we argue that these drawbacks potentially both stem from the same underlying issue — low mutual information between the model’s input features and task-specific output during optimization. This low mutual information arises from constraining both tasks within a shared parameter space in the shared encoder, as depicted in the parameter space analysis in Figure 1(a). This constraint limits the representation capability of the shared encoders, resulting in gradient conflict. Thus, it necessitates the manual design of task-specific heads to compensate the potential performance degradation caused by gradient conflicts.
@@
To address the above limitations, we aim to increase the mutual information between the model’s input features and the task-specific outputs. Partitioning the parameter space into distinct subspaces has proven to be effective for enhancing mutual information, particularly when addressing diverse tasks (Ueda and Nishikawa, 2020). It reduces task interference and enables the model to learn task-specific predictions through different task optimization subspaces. To naturally achieve such partitioning, we design a generative paradigm that leverages instruction tuning in pre-trained generative models. Instruction tuning partitions the parameter space by introducing distinct prompt for each task, effectively tailoring a unique optimization subspace for S&R as illustrated in Figure 1(b).
@@
In summary, the main contributions are threefold: (1) We propose GenSR, a generative paradigm designed to unify S&R tasks through dual representation learning and S&R task unifying. (2) We introduce a novel perspective for comparing discriminative and generative paradigms based on mutual information. (3) We conduct comprehensive experiments on two public datasets, including S&R performance analysis, mutual information estimation, optimization space visualization, and ablation study of different modules, to validate GenSR’s effectiveness and efficiency compared to other baselines across both S&R tasks.
@@ 2. Related Work
Unifying search and recommendation. Unifying S&R by leveraging data from both tasks can enhance user modeling and item understanding, broadly categorized into two approaches: 1) Search-enhanced recommendation (Wu et al., 2019), which uses user search behavior as supplementary to uses’ representation, thus assisting recommendation. SESRec (Si et al., 2023) utilizes contrastive learning to extract similar and dissimilar user interests from S&R, and then leverages users’ search interests for recommendation. 2) Unifying S&R (Yao et al., 2021), which shares S&R data and utilizes one model to boost the performance on both two tasks. For example, UniSAR (Shi et al., 2024) models fine-grained user behavior transitions between S&R through extraction, alignment, and fusion. However, they follow the discriminative paradigm, which introduces gradient conflict and manual design complexity. Recently, Penha et al. (2024) proposed a generative model for unifying S&R. While promising, it treats S&R data independently, ignoring their joint interaction history—a rich source of contextual signals for user modeling.
In contrast, we introduce a generative paradigm that enhances mutual information and jointly leverages S&R interactions to better model user and item representations.
Generative recommendation. Pre-trained generative models have garnered attention in recommendation for their contextual understanding and global knowledge (Lin et al., 2024). Early works like P5 (Geng et al., 2022), TALLRec (Bao et al., 2023b), and M6-Rec (Cui et al., 2022) fine-tune these models with prompts and recommendation data. Later methods such as BIGRec (Bao et al., 2023a), IDGenRec (Tan et al., 2024), and TIGER (Rajput et al., 2023) enhance grounding in item space and semantic relevance. Recent efforts further highlight the critical role of CF information in generative recommendation, as CF plays a crucial role in understanding user and item representations in recommendation systems (Wang et al., 2024; Kim et al., 2024). However, integrating search data into generative recommendations remains underexplored and holds potential for capturing users’ immediate intent to boost personalization.
Personalized search. Personalized search aims to deliver search results tailored to users’ preferences, behaviors, and contextual information. For instance, HEM learns user representations from user profiles to provide personalized search rankings (Ai et al., 2017). TEM enhances this by incorporating user search histories, allowing it to better capture users’ preferences and historical behaviors (Bi et al., 2020).
@@
Generative model for ranking. Due to the advanced contextual understanding capabilities of pre-trained language models, generative models for ranking have emerged as a novel approach to directly model the relevance between queries and items by framing ranking as a sequence generation task. For example,  Nogueira et al. (2020) adapts T5 for document ranking by training it to generate relevance labels (“true” or “false”) for query-document pairs. Similarly,  Ferraretto et al. (2023) introduces a generative ranking model trained to output both relevance labels and explanations for query-document pairs. Despite these advances, it’s promising to incorpoate recommendation data to extract users’ preferences which helps to boost the performance on the search side.
@@ 3. Task Formulation
We formulate the task of unifying S&R under both discriminative and generative paradigms. The objective is to integrate users’ historical S&R interactions into a single model capable of handling both tasks based on user needs. Specifically, let $\mathcal{U}$ and $\mathcal{I}$ be the sets of users and items. For a user $u\in\mathcal{U}$, the model predicts preference for a candidate item $i_{c}$ in either a search or recommendation context. To achieve this, the model takes as input three types of information for each user $u$: (1) Interaction history $H_{u}=\{(i_{1},b_{1}),(i_{2},b_{2}),\cdots,(i_{N},b_{N})\}$, where $i_{n}\in\mathcal{I}$ is the $n-$th item in the interaction history, $N$ is the total number of interactions, $b_{n}$ is the type of behavior for $n-$th item, with ”src” indicating search and ”rec” indicating recommendation. (2) Candidate item $i_{c}$, which is the item candidate pool used for evaluating whether the user $u$ will like (i.e., click) this item. (3) Query $q$, issued if the current task is search; otherwise, $q$ is empty for recommendation.
@@
(1) $\mathop{\min}_{\Phi}\{\mathcal{L}_{\Phi}=\mathcal{L}{s}+\gamma\mathcal{L}{r}\},$
@@
Generative paradigm. Differently, in the generative paradigm, the input information is transformed into a unified structured prompt $x$, which is tokenized before being fed into the model for instruction tuning. The model generates “Yes” or “No” to indicate user preference via constrained generation.
The training objective minimizes the negative the negative log-likelihood of the users’ preference $y$ conditioned on input $x$:
@@ 4. Method
Figure 2. GenSR structure. Given S&R history, GenSR uses dual representation learning to obtain CF and semantic embeddings, and unifies both tasks via contrastive learning combined with instruction tuning to generate task-specific outputs.
We propose GenSR, a novel generative paradigm for unifying search and recommendation (S&R). GenSR has two modules: 1) Dual Representation learning, which independently learns item representations from both the CF view and the semantic view. 2) S&R Task Unifying, which utilizes contrastive learning together with instruction tuning to generate task-specific outputs effectively. The overall architecture of GenSR is shown in Figure 2.
@@ 4.1.1. Dual-view representation
From this history, we compute the corresponding CF representations $E_{u}^{\text{cf}}=\{e_{1},e_{2},\cdots,e_{N}\}^{T}\in\mathbb{R}^{N\times d}$ using a pre-trained CF embedding table ||×$\mathcal{E}\in\mathbb{R}^{|\mathcal{I}|\times d}$. Here, each $e_{i}$ is a $d$-dimensional embedding, $|\mathcal{I}|$ denotes the total number of items in the corpus, and $d$ represents the dimensionality of the CF embeddings. [^1] The CF embedding table $\mathcal{E}$ is derived from training LightGCN using only the recommendation behaviors in the dataset (He et al., 2020).

### v3

# Unifying Search and Recommendation with Dual-View Representation Learning in a Generative Paradigm

@@ Abstract.
Recommender systems and search engines are critical components of online platforms, with recommender systems proactively providing information to users, whereas search engines enable users to seek information actively. Unifying both tasks in a shared model is promising since it can enhance user modeling and item understanding. Previous approaches mainly follow a discriminative paradigm, utilizing shared encoders to process input features and task-specific heads to perform each task. However, this paradigm faces two key challenges: gradient conflict and manual architecture design overhead. From the information theory perspective, these challenges potentially both stem from the same issue — low mutual information between input features and task-specific output during optimization.
To address this, we propose GenSR, a novel generative paradigm for unifying search and recommendation (S&R), which leverages task-specific prompts to partition the model’s parameter space into subspaces, thereby enhancing mutual information. To construct effective subspaces for each task, GenSR first prepares informative representations for each subspace and then optimizes both subspaces in one unified model. Specifically, GenSR consists of two main modules: (1) Dual Representation Learning, which independently models collaborative and semantic historical information to derive expressive item representations; and (2) S&R Task Unifying, which utilizes contrastive learning together with instruction tuning to generate task-specific outputs effectively. Extensive experiments on two benchmarks show GenSR outperforms state-of-the-art methods across S&R tasks.
Unifying Search and Recommendation, Generative Recommendation, Multi-task Learning
@@ 1. Introduction
Recommender systems and search engines have become indispensable components of modern online service platforms, including e-commerce websites and social media networks (Zhang et al., 2024; Wu et al., 2024). Usually, search and recommendation (S&R) tasks are trained on separate models (Bhattacharya et al., 2024). Recently, advances in information retrieval show that unifying these two tasks within a shared model offers significant advantages (Yao et al., 2021; Si et al., 2023): First, joint modeling of user behaviors across both S&R provides a more comprehensive understanding of user preferences. Second, leveraging interactions from both task domains enriches item-side interactions, enhancing the understanding and delivery of items.
@@
Figure 1. (a) and (b) illustrate the discriminative and generative paradigms for unifying S&R, respectively, highlighting their differences in parameter space. While the discriminative paradigm results in low mutual information by optimizing S&R tasks within a shared parameter space, the generative paradigm enhances mutual information by partitioning the space into distinct subspaces through instructions.
@@
Existing efforts to unify S&R mostly follow a discriminative paradigm, which utilizes shared encoders to process input features (e.g., user interaction history) and task-specific heads to predict whether a user will click on an item in S&R tasks, as illustrated in Figure 1(a) (Shi et al., 2024; Xie et al., 2024). However, such paradigm has two drawbacks: 1) Gradient conflict: The differing optimization objectives of S&R, where search primarily focuses on modeling query-item relevance based on the current query context, while recommendation emphasizes capturing user preferences by analyzing user’s historical interactions, lead to gradient conflicts in the shared encoder (Shi et al., 2023; He et al., 2024). This often results in negative cosine similarity between the task gradients and interferes with effective model optimization (Yu et al., 2020) (see empirical evidence in Section 6.5). 2) Manual architecture design overhead: The design of task-specific heads requires substantial manual effort, increasing system complexity and reducing scalability and extensibility (Rahimian et al., 2023; Xie et al., 2024).
@@
Inspired by information theory, we argue that these drawbacks potentially both stem from the same underlying issue — low mutual information between the model’s input features and task-specific output during optimization. This low mutual information arises from constraining both tasks within a shared parameter space in the shared encoder, as depicted in the parameter space analysis in Figure 1(a). This constraint limits the representation capability of the shared encoder, resulting in gradient conflict. Consequently, it necessitates manual designs of task-specific heads to compensate the potential performance degradation caused by gradient conflicts.
@@
To address the above limitations, we aim to increase the mutual information between the model’s input features and the task-specific outputs. Partitioning the parameter space into distinct subspaces has proven to be effective for enhancing mutual information, particularly when addressing diverse tasks (Ueda and Nishikawa, 2020). It reduces task interference and enables the model to learn task-specific predictions through different task optimization subspaces. To naturally achieve such partitioning, we design a generative paradigm that leverages instruction tuning in pre-trained generative models. Instruction tuning partitions the parameter space by introducing distinct prompts for each task, effectively tailoring a unique optimization subspace for S&R as illustrated in Figure 1(b).
@@
In summary, our main contributions are as follows:
  1. (1)
    We propose GenSR, a generative paradigm designed to unify S&R tasks through dual representation learning and S&R task unifying.
  2. (2)
    We introduce a novel perspective for comparing discriminative and generative paradigms based on mutual information.
  3. (3)
    We conduct comprehensive experiments on two public datasets, including S&R performance analysis, mutual information estimation, parameter space visualization, gradient conflict analysis and ablation study of different modules, to validate GenSR’s effectiveness and efficiency compared to other baselines across both S&R tasks.
@@ 2. Related Work
2.1. Unifying search and recommendation
Recent research highlights that search and recommendation (S&R) behaviors are inherently complementary, as each provides unique signals that enhance user modeling and item understanding (Penha et al., 2024; Shi et al., 2025). By jointly leveraging data from both tasks, it is possible to achieve more accurate intent modeling and representation learning, thereby improving the performance of both tasks. As a result, an increasing number of studies have focused on unifying S&R. These efforts can be broadly categorized into two main directions:
(1) Search-enhanced recommendation (Wu et al., 2019), which integrates search behavior into the recommendation scenario to improve user preference modeling. This line of research assumes that users’ search queries convey semantic intent signals that are often missing from the behavioral data used in recommendation systems. For instance, SESRec (Si et al., 2023) utilizes contrastive learning to extract similar and dissimilar user interests from S&R, and then leverages users’ search interests for recommendation. Query-SeqRec (He et al., 2022) effectively incorporates explicit user queries from the search side, alongside item interactions, for improved user intent modeling and recommendation performance.
(2) Unifying S&R (Yao et al., 2021), which shares S&R data and utilizes one model to boost the performance on both tasks. This line of work typically utilizes shared representations or architectures to model S&R behaviors within a unified framework. For example, UnifiedSSR (Xie et al., 2024) jointly models user behavior history in S&R scenarios using a parameter-sharing dual-branch network and an intent-oriented session module. UniSAR (Shi et al., 2024) models fine-grained user behavior transitions between S&R through extraction, alignment, and fusion.
@@ 2.1. Unifying search and recommendation
However, they follow the discriminative paradigm, which introduces gradient conflict and manual design complexity. Recently, Penha et al. (2024) proposed a generative model for unifying S&R. While promising, it treats S&R data independently, ignoring their joint interaction history—a rich source of contextual signals for user modeling. In contrast, we introduce a generative paradigm that jointly leverages S&R interactions to better model user and item representations and enhances mutual information.
2.2. Generative recommendation
Generative models, notably Generative Adversarial Networks (GANs) (Wang et al., 2017; Gao et al., 2021; Jin et al., 2020) and Variational Autoencoders (VAEs) (Ma et al., 2019; Liang et al., 2018; Zhang et al., 2017), are pivotal for personalized recommendations but face structural limitations (Sohl-Dickstein et al., 2015; Kingma et al., 2016). Nowadays, pre-trained generative models have garnered attention in recommendation for their contextual understanding and global knowledge (Lin et al., 2024), which better capture user preferences from interactions. Early studies focused on fine-tuning these models with tailored prompts and recommendation data, e.g., P5 (Geng et al., 2022), TALLRec (Bao et al., 2023b), and M6-Rec (Cui et al., 2022). For instance, TALLRec introduces an effective and efficient instruction tuning framework to align large language models (LLMs) with recommendation tasks, addressing the challenge that LLMs are not originally pre-trained for such purposes. Subsequent advancements, notably BIGRec (Bao et al., 2023a), IDGenRec (Tan et al., 2024) and TIGER (Rajput et al., 2023), extend this approach by grounding outputs in specific item spaces or incorporating semantic details to refine recommendations. For example, BIGRec first fine-tunes LLMs to generate meaningful item descriptions and then maps these to actual items, enhancing recommendation performance. IDGenRec proposes learning unique, semantically rich textual IDs for items to seamlessly integrate them into LLM-based generative recommender systems. Recently, researchers have recognized the critical role of CF information in generative recommendation, as CF plays a crucial role in understanding user and item representations in recommendation systems (Wang et al., 2024; Kim et al., 2024).
However, integrating search data into generative recommendations remains underexplored and holds potential for capturing users’ immediate intent to boost personalization.
2.3. Personalized search
Personalized search aims to deliver items tailored to users’ preferences, behaviors, and contextual information. Early methods learn user representations from profiles or search histories to re-rank items based on personalized relevance. For instance, HEM learns user representations from user profiles to provide personalized search rankings (Ai et al., 2017). TEM enhances this by incorporating user search histories, allowing it to better capture users’ preferences and historical behaviors (Bi et al., 2020). Beyond these, some approaches tackle key challenges such as sparse user behavior and noisy interaction data. For instance, CoPPS (Dai et al., 2023) leverages contrastive learning to enhance user representations under sparse and noisy search scenarios. QIN (Guo et al., 2023) incorporates cascaded filtering to extract query-relevant behaviors and employs fused attention to distill salient signals from long-tail history. However, most personalized search methods rely solely on search-side signals, overlooking the rich collaborative filtering information available in recommendation interactions. This limits their ability to effectively model long-term user preferences.
In contrast, our proposed model adopts a unified framework that integrates both recommendation and search behaviors within a generative paradigm, enabling the construction of task-aware prompts and richer user representations to improve retrieval quality.
2.4. Generative model for ranking
@@
Due to the advanced contextual understanding capabilities of pre-trained language models, generative models for ranking have emerged as a novel approach to directly model the relevance between queries and items by casting ranking as a sequence generation problem. These methods approach ranking by generating sequences conditioned on input queries and documents, shifting away from traditional scoring-based frameworks. By treating ranking as a text generation task, they are capable of modeling complex interactions between queries and items in a more flexible and contextualized manner. For example, Nogueira et al. (2020) adapts T5 for document ranking by training it to generate relevance labels (“true” or “false”) for query-document pairs. Similarly, Ferraretto et al. (2023) introduces a generative ranking model trained to output both relevance labels and explanations for query-document pairs.
Despite these advances, it’s promising to incorporate recommendation data to extract users’ preferences which helps to boost the performance on the search side.
@@ 3. Task Formulation
We formulate the task of unifying S&R within both discriminative and generative paradigms. The goal is to integrate users’ historical S&R interactions into a single model capable of performing both tasks according to user needs. In the search setting, the objective is to retrieve documents relevant to a user’s query while leveraging their historical S&R interactions. In the recommendation setting, the objective is to suggest items based on the user’s historical S&R interactions. Formally, let $\mathcal{U}$ and $\mathcal{I}$ be the sets of users and items. For a user $u\in\mathcal{U}$, the model predicts preference for a candidate item $i_{c}$ in either a search or recommendation context. To achieve this, the model takes as input three types of information for each user $u$:
  1. (1)
    Interaction history $H_{u}=\{(i_{1},b_{1}),(i_{2},b_{2}),\cdots,(i_{N},b_{N})\}$, where $i_{n}\in\mathcal{I}$ is the $n$-th item in the interaction history, $N$ is the total number of interactions, $b_{n}$ is the type of behavior for $n$-th item, with ”src” indicating search and ”rec” indicating recommendation.
  2. (2)
    Candidate item $i_{c}$, which is the item candidate pool used for evaluating whether the user $u$ will like (i.e., click) this item.
  3. (3)
    Query $q$, issued if the current task is search; otherwise, $q$ is empty for recommendation.
@@
(1) $\mathop{\min}_{\Phi}\{\mathcal{L}_{\Phi}=\mathcal{L}{s}(H_{u},i_{c},q)+\gamma\mathcal{L}{r}(H_{u},i_{c})\},$
@@
Generative paradigm. Differently, in the generative paradigm, we cast both S&R as conditional text generation. The input information is transformed into a unified structured prompt $x$, which is tokenized before being fed into the model for instruction tuning. The model generates “Yes” or “No” to indicate user preference via constrained generation.
The training objective minimizes the negative log-likelihood of the users’ preference $y$ conditioned on input $x$ in an auto-regressive manner:
@@
4. Mutual Information Analysis of GenSR
In this section, we conduct an in-depth analysis of mutual information to support our claim that the proposed generative paradigm effectively enhances mutual information compared to traditional discriminative methods. We begin by defining mutual information from an information-theoretic perspective, and then present the theoretical evidence to substantiate this claim.
4.1. Information-theoretic perspective
From an information-theoretic standpoint, the representational capability of a unified S&R model can be examined via the mutual information between the input $X$ and the output $Y_{t}$ for a task $t\in\{S,R\}$. Mutual information, denoted as $I(X;Y_{t})$, measures how much information the input $X$ provides about the output $Y_{t}$, quantifying their dependence. By definition:
(3) $I(X;Y_{t})\;=\;H(Y_{t})\;-\;H(Y_{t}\mid X),$
where $H(Y_{t})$ is the entropy of $Y_{t}$, and $H(Y_{t}\mid X)$ is the conditional entropy of $Y_{t}$ given $X$.
A higher mutual information indicates that the input $X$ contains more predictive information about the output $Y_{t}$, reflecting a stronger alignment between inputs and task-specific outputs, i.e., higher model representation capability. Conversely, lower mutual information implies that predicted output fail to fully capture the distribution and pattern of input, reflecting a lower model representation capability. Therefore, mutual information $I(X;Y_{t})$ serves as a good indicator for model representation capability (Bachman et al., 2019).
4.2. Theoretical analysis
In this section, we aim to show the effectiveness of our model GenSR, based on the aforementioned mutual information. Specifically, we conduct a theoretical analysis to show this utilizing the following theorem:
Theorem 4.1.
Suppose the input distribution $X\sim\mathcal{N}(\mu,\Sigma)$, and the noise $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}I)$ is independent of $X$. Consider two paradigms: generative (with parameters $\phi$) and discriminative (with parameter $\theta$). Assume the loss functions $\mathcal{L}_{\theta}$ and $\mathcal{L}_{\phi}$ admit second-order Taylor expansions around their respective true parameters with bounded remainder terms, [^1] This assumption is common in statistical learning and information retrieval (Prasad et al., 2017; Kishida, 2003). then the mutual information between the input and output satisfies:
We provide a detailed proof of Theorem 4.1 from the view of parameter approximation separability in the Appendix. Intuitively, we prove that generative paradigm can enhance mutual information in S&R scenario compared with discriminative paradigm under certain input distribution assumption.
@@ 5. Method
Figure 2. GenSR structure. Given S&R history, GenSR first utilizes dual representation learning to obtain CF and semantic history embeddings. The S&R task unifying module then integrates these dual representations with task-specific prompts into a shared generative model, leveraging contrastive learning for view alignment and instruction tuning to effectively generate task-specific outputs for both search and recommendation.
In this section, we introduce GenSR, a novel generative paradigm for unifying search and recommendation (S&R). GenSR has two modules: 1) Dual Representation learning, which independently learns item representations from both the CF view and the semantic view. 2) S&R Task Unifying, which utilizes contrastive learning together with instruction tuning to generate task-specific outputs effectively. The overall architecture of GenSR is shown in Figure 2.
@@ 5.1.1. Dual-view representation
From this history, we compute the corresponding CF representations $E_{u}^{\text{cf}}=\{e_{1},e_{2},\cdots,e_{N}\}^{T}\in\mathbb{R}^{N\times d}$ using a pre-trained CF embedding table $\mathcal{E}\in\mathbb{R}^{|\mathcal{I}|\times d}$. Here, each $e_{i}$ is a $d$-dimensional embedding, $|\mathcal{I}|$ denotes the total number of items in the corpus, and $d$ represents the dimensionality of the CF embeddings. [^2] The CF embedding table $\mathcal{E}$ is derived from training LightGCN using only the recommendation behaviors in the dataset (He et al., 2020).